OpenStudy (anonymous):

2. Saturn has a radius of about 9.0 earth radii, and a mass 95 times the Earth’s mass. Estimate the gravitational field on the surface of Saturn compared to that on the Earth

OpenStudy (anonymous):

@Abhisar

OpenStudy (anonymous):

@pooja195

OpenStudy (abhisar):

Do you know the formula we use to calculate gravitational force?

OpenStudy (anonymous):

F = G \frac{m_1 m_2}{r^2}\ ,

OpenStudy (abhisar):

Yes, Now for field we just replace the other mass with 1 and R with radius So the formula for field would be \(\sf \huge \frac{G\times M}{R^2}\) Let the mass of earth be M and radius be R \(\therefore\) Field will be \(\sf \huge \frac{G\times M}{R^2}\). Let this be equal to a random quantity F. In next step calculate the field of Saturn by replacing M with 95M and R with9R. Compare the result with F

OpenStudy (anonymous):

\[\frac{ G*95M }{9R^2 }\]

OpenStudy (abhisar):

You missed it, The denominator will be \(\sf (9R)^2\) = ?

OpenStudy (anonymous):

?

OpenStudy (abhisar):

\[\sf \frac{ G*95M }{(9R)^2 }\]

OpenStudy (abhisar):

Getting it....because you replace R with (9R)

OpenStudy (anonymous):

G∗95M\[G∗95M(9R)2\] (9R)2

OpenStudy (anonymous):

so the answer is (9R)^2?

OpenStudy (anonymous):

@Abhisar

OpenStudy (abhisar):

No. You need to compare \[\sf \frac{ G*95M }{(9R)^2 }\] with Field of earth we calculated earlier.

OpenStudy (anonymous):

\[11.49m/s^2\]